### v2

# Minimal Cellular Resolutions of Path Ideals

Abstract.
In this paper, we prove that the path ideals of both paths and cycles have minimal cellular resolutions. Specifically, these minimal free resolutions coincide with the Barile-Macchia resolutions for paths, and their generalized counterparts for cycles. Furthermore, we identify edge ideals of cycles as a class of ideals that lack a minimal Barile-Macchia resolution, yet have a minimal generalized Barile-Macchia resolution.
1. Introduction
It has been a powerful approach to associate a combinatorial object with an algebraic one and study its algebraic properties via combinatorics [12, 13, 16, 17, 18]. In the spirit of this approach, the algebraic objects of interest in this paper are path ideals, while the combinatorial counterparts are graphs. Specifically, our focus is on studying the path ideals of graphs and their minimal free resolutions by leveraging the underlying structure of the graphs. Central to this work is the use of (generalized) Barile-Macchia resolutions from [9]. Such resolutions fall under the umbrella of Morse resolutions, a class of cellular resolutions first introduced by Batzies and Welker in [5]. These resolutions are obtained using homogeneous acyclic matchings, a concept from discrete Morse theory. In addition to the Barile-Macchia resolution, other examples of Morse resolutions have been introduced in the literature. One recent example is the pruned resolutions from [3].
[…]
In this paper, our goal is to construct minimal resolutions for path ideals of paths and cycles. While there is literature discussing and providing explicit formulas for the (graded) Betti numbers of these ideals [2, 14], no construction has yet been provided for their minimal resolution. We achieve this by working with (generalized) Barile-Macchia resolutions from [9], thereby expanding the class of ideals for which these resolutions are minimal.
Our two main results are:
  1. (1)
    We establish that path ideals of paths have the bridge-friendly property (Theorem 3.8). This property ensures the minimality of Barile-Macchia resolutions as described in [9]. Thus, we determine that the path ideals of paths admit a minimal Barile-Macchia resolution. From this, we derive formulas for their projective dimension and graded Betti numbers, recovering results from [2] and [7] (Corollary 3.15, Theorem 3.16).
  2. (2)
    We shift our focus to path ideals of cycles. In this context, Barile-Macchia resolutions are not always minimal. One instance is the edge ideal of a 9-cycle, $I_{2}(C_{9})$, which, as indicated in [9], does not have any minimal Barile-Macchia resolution. Nonetheless, we prove that path ideals of cycles have a minimal generalized Barile-Macchia resolution (Theorem 4.16).
Our results on minimal cellular resolutions of path ideals for paths and cycles generalize the results from [3], where it is shown that edge ideals of paths and cycles have minimal pruned resolutions.
Both Barile-Macchia and generalized Barile-Macchia resolutions are induced by homogeneous acyclic matchings, called Barile-Macchia and generalized Barile-Macchia matchings, respectively. At the heart of the Barile-Macchia matching construction is the comparison of least common multiples of subsets of the minimal generating set ${\mathcal{G}}(I)$ of a monomial ideal $I$, with respect to a total order on ${\mathcal{G}}(I)$. An algorithm for producing Barile-Macchia matchings was introduced by the first two authors in [9]. In addition, MorseResolutions Macaulay2 package dedicated to Barile-Macchia resolutions were introduced by the first two authors and O’Keefe in [10]. While the generalized version adopts the same foundational principle, it extends to multiple total orders on ${\mathcal{G}}(I)$ as discussed in Section 4. For further details, refer to [9].
An important concept in relating homogeneous acyclic matchings and free resolutions is critical cells: These are subsets of ${\mathcal{G}}(I)$ that remain untouched by a homogeneous acyclic matching of $I$. In [5], it was shown that these cells are in one-to-one correspondence with the ranks of free modules from (generalized) Barile-Macchia resolutions. Thus, in this paper, we focus on characterizing critical cells of path ideals using bridges, gaps and true gaps – simple yet powerful concepts rooted in the graph’s structure as introduced in [9]. We use these notions to produce minimal free resolutions of path ideals of paths and cycles.
A key observation concerning the critical cells of path ideals of paths is that two distinct critical cells have different least common multiples. This observation proves useful in constructing a critical cell of maximum size, which in turn allows us to deduce a formula for the projective dimension of path ideals of paths. Additionally, this insight is helpful in identifying all critical cells of path ideals of cycles. Such cells consists of critical cells of path ideals of induced paths as well as critical cells whose least common multiple has the largest multidegree – for a cycle on $n$ vertices, it would be $x_{1}\dots x_{n}$. Consequently, in the paper’s final section, our study primarily focuses on identifying the critical cells of the latter type for cycles.
The structure of this paper is as follows. In Section 2, we revisit essential concepts and results relevant to Morse resolutions, as well as Barile-Macchia matchings and resolutions. In Section 3, we explore path ideals of paths. We start this section by delving into the characterizations of bridges, gaps, and true gaps specific to paths. With these characterizations in hand, we obtain the bridge-friendliness and affirm the minimality of the associated Barile-Macchia resolution. This paves the way for us to introduce formulas for the projective dimension and to provide a recursive formula for graded Betti numbers. Lastly, in Section 4, we turn our attention to path ideals of cycles, offering both a review of and insights into the application of generalized Barile-Macchia resolutions. After characterizing the bridges, gaps, and true gaps, we verify the minimality of their generalized Barile-Macchia resolutions by drawing upon our earlier findings on path ideals.
2. Preliminaries
2.1. Morse resolutions
[…]
$E(G_{X}^{A})=(E\setminus A)\cup\{(\sigma^{\prime},\sigma)\mid(\sigma,\sigma^{\prime})\in A\}.$
Essentially, $G_{X}^{A}$ is derived from $G_{X}$ by reversing the direction of edges belonging to $A$.
Central to our discussion on a Morse resolution of $I$ is the notion of homogeneous acyclic matchings from discrete Morse theory, a concept we revisit below.
Definition 2.1.
A subset $A\subseteq E$ is called a homogeneous acyclic matching of $I$ if the following conditions hold:
  1. (1)
    (matching) Each cell appears in, at most, one edge of $A$.
  2. (2)
    (acyclicity) The graph $G^{A}_{X}$ does not contain a directed cycle.
  3. (3)
    (homogeneity) For any edge $(\sigma,\sigma^{\prime})$ in $A$, we have $\operatorname{lcm}(\sigma)=\operatorname{lcm}(\sigma^{\prime})$.
A cell $\sigma$ that does not appear in any edge of $A$ is called an $A$-critical cell of $I$. In contexts where there is no ambiguity, we simply refer to it as critical.
Recall that $X$ is a $\mathbb{Z}^{N}$-graded complex that induces a free resolution $\mathcal{F}$ where $\mathcal{F}_{r}$ is the free $R$-module with a basis indexed by all cells of cardinality $r$. The differentials $\partial_{r}\colon\mathcal{F}_{r}\to\mathcal{F}_{r-1},$ are defined as
$\partial_{r}(\sigma)=\sum_{\begin{subarray}{c}\sigma^{\prime}\subseteq\sigma,\\ |\sigma^{\prime}|=r-1\end{subarray}}[\sigma:\sigma^{\prime}]\frac{\operatorname{lcm}(\sigma)}{\operatorname{lcm}(\sigma^{\prime})}\sigma^{\prime}.$
where $[\sigma:\sigma^{\prime}]$ denotes the coefficient of $\sigma^{\prime}$ in the boundary of $\sigma$ and is either $1$ or $-1$. The complex $\mathcal{F}$ is called the Taylor resolution of $R/I$.
Morse resolutions are refinements of the Taylor resolution. Each homogeneous acyclic matching yields a Morse resolution, and these resolutions may coincide. To precisely define the differentials of a Morse resolution, we need to introduce some additional terminology.
Given a directed edge $(\sigma,\sigma^{\prime})\in E(G^{A})$, set
$m(\sigma,\sigma^{\prime})=\begin{cases}-[\sigma^{\prime}:\sigma]&\text{if }(\sigma^{\prime},\sigma)\in A,\\ [\sigma:\sigma^{\prime}]&\text{otherwise}.\end{cases}$
A gradient path $\mathcal{P}$ from $\sigma_{1}$ to $\sigma_{t}$ is a directed path
$\mathcal{P}\colon\sigma_{1}\to\sigma_{2}\to\cdots\to\sigma_{t}$
in $G^{A}_{X}$. Similarly, set $m(\mathcal{P})=m(\sigma_{1},\sigma_{2})\cdot m(\sigma_{2},\sigma_{3})\cdots m(\sigma_{t-1},\sigma_{t}).$
We are now ready to recall Morse resolutions of monomial ideals.
Theorem 2.2.
[5, Proposition 2.2, Proposition 3.1, Lemma 7.7] Let $A$ be a homogeneous acyclic matching of $I$. Then $A$ induces a cellular resolution $\mathcal{F}_{A}$, where:
  • $(\mathcal{F}_{A})_{r}$ is the free $R$-module with a basis indexed by all critical cells of cardinality $r$.
  • The differentials $\partial_{r}^{A}:(\mathcal{F}_{A})_{r}\to(\mathcal{F}_{A})_{r-1}$ are defined by:
[…]
    from
    σ to σ′′
[…]
The resulting (cellular) free resolution $\mathcal{F}_{A}$ is called the Morse resolution of $R/I$ associated to $A$. Furthermore, $\mathcal{F}_{A}$ is minimal if for any two $A$-critical cells $\sigma$ and $\sigma^{\prime}$ with $|\sigma^{\prime}|=|\sigma|-1$, we have $\operatorname{lcm}(\sigma)\neq\operatorname{lcm}(\sigma^{\prime})$.
2.2. Barile-Macchia matchings and Barile-Macchia resolutions.
In this subsection, we recall Barile-Macchia matchings and resolutions. To ensure that this paper is self-contained and accessible, we present relevant definitions and results from [9] that are instrumental to our discussions.
Given Theorem 2.2, the key to producing a Morse resolution of $R/I$ is finding a homogeneous acyclic matching of $I$. However, systematically crafting such a matching for any monomial ideal is a challenging task. A recent development in this direction is the Barile-Macchia algorithm, as presented in [9, Theorem 2.8], which produces a homogeneous acyclic matching. In the context of [9], a matching arising from this algorithm is called a Barile-Macchia matching, and the resulting resolution is called a Barile-Macchia resolution.
Below, we recall some of the terminology useful for discussing Barile-Macchia resolutions. Throughout this section, we fix a total order $(\succ_{I})$ on ${\mathcal{G}}(I)$. For simplicity, we write $S\setminus s$ (resp, $S\cup s$) instead of $S\setminus\{s\}$ (resp, $S\cup\{s\}$) where $S$ is a set and $s\in S$ (resp, $s\notin S$).
Additionally, throughout our discussion, we use the terms “subsets of ${\mathcal{G}}(I)$” and “cells of $I$” interchangeably. By cells of $I$, we refer to cells of the corresponding Taylor simplex $X$.
Definition 2.3.
Let $\sigma$ be a subset of ${\mathcal{G}}(I)$. A monomial $m$ is called a bridge of $\sigma$ if $m\in\sigma$ and removing $m$ from $\sigma$ does not change the $\operatorname{lcm}$, i.e., $\operatorname{lcm}(\sigma\setminus m)=\operatorname{lcm}(\sigma)$.
Notation 2.4.
If $\sigma$ has a bridge, the notation $\operatorname{sb}(\sigma)$ denotes the smallest bridge of $\sigma$ with respect to $(\succ_{I})$. We set $\operatorname{sb}(\sigma)=\emptyset$ if $\sigma$ has no bridges.
To fully understand the terms and context we discuss, in Algorithm 1 we recall the Barile-Macchia algorithm as outlined in [9, Algorithm 2.9]. The Barile-Macchia Algorithm systematically constructs a matching in $G_{X}$ based on the concept of bridges within subsets on ${\mathcal{G}}(I)$ and a fixed total order $(\succ_{I})$ of ${\mathcal{G}}(I)$. As it was shown in [9, Theorem 2.11], this process always produces a homogeneous acyclic matching.
Algorithm 1 Barile-Macchia Algorithm
Input: A total order $(\succ_{I})$ on ${\mathcal{G}}(I)$.
Output: Set of directed edges $A$ in $G_{X}$.
1:$A\leftarrow\emptyset$
2:Ω{all subsets of $\Omega\leftarrow\{\text{all subsets of }{\mathcal{G}}(I)\text{ with cardinality at least }3\}$
3:while $\Omega\neq\emptyset$ do
4:     Pick $\sigma\in\Omega$ with maximal cardinality
5:     Remove $\{\sigma,\sigma\setminus\operatorname{sb}(\sigma)\}$ from $\Omega$
6:     if $\operatorname{sb}(\sigma)\neq\emptyset$ then
7:         Add edge $(\sigma,\sigma\setminus\operatorname{sb}(\sigma))$ to $A$
8:for all distinct edges $(\sigma,\tau)$ and $(\sigma^{\prime},\tau^{\prime})$ in $A$ with $\tau=\tau^{\prime}$ do
9:     if $\operatorname{sb}(\sigma^{\prime})\succ_{I}\operatorname{sb}(\sigma)$ then
10:         Remove $(\sigma^{\prime},\tau^{\prime})$ from $A$
11:     else
12:         Remove $(\sigma,\tau)$ from $A$      return $A$
Definition 2.5.
Given a Barile-Macchia matching $A$ of $I$:
  1. (1)
    For any edge $(\sigma,\tau)$ in the final $A$ from Algorithm 1 , the cell $\sigma$ is called type-2 while the cell $\tau$ is called type-1.
  2. (2)
    During the execution of Algorithm 1, if a directed edge $(\sigma,\tau)$ is added to $A$, the cell $\sigma$ is called potentially-type-2. It is important to note that this edge may not persist in the final $A$ produced by Algorithm 1.
Remark 2.6.
If a subset of ${\mathcal{G}}(I)$ has a bridge, then it is either type-1 or potentially-type-2.
In earlier discussions, we emphasized the one-to-one correspondence between the ranks of the free $R$-modules in a Barile-Macchia resolution of $R/I$ and the $A$-critical cells of $I$. Here, $A$ represents the Barile-Macchia matching of $I$ with respect to $(\succ_{I})$. Thus, delving into the nature of $A$-critical cells, along with the remaining cells, is crucial for a deeper comprehension of the Barile-Macchia resolution of $R/I$. It is important to point out that critical cells of $I$ consists of those cells either left out of $A$ during the Barile-Macchia Algorithm or added initially but later excluded in the final refinement of the algorithm. We name these critical cells as follows:
Definition 2.7.
Let $\sigma$ be a subset of ${\mathcal{G}}(I)$. If $\sigma$ is never added to $A$ in any of the steps throughout Algorithm 1, it is called absolutely critical. If $\sigma$ is potentially-type-2 but not type-2 (initially added to $A$ but removed in the final $A$ produced by Algorithm 1), we call it fortunately critical.
The following concepts from [9] are useful in characterizing critical and non-critical cells of $I$.
Definition 2.8.
[9, Definition 2.19] Let $m,m^{\prime}\in{\mathcal{G}}(I)$ and $\sigma$ be a subset of ${\mathcal{G}}(I)$.
  1. (1)
    We say that $m$ dominates $m^{\prime}$ if and only if $m\succ_{I}m^{\prime}$.
  2. (2)
    The monomial $m$ is called a gap of $\sigma$ if $m\notin\sigma$, and $\operatorname{lcm}(\sigma\cup m)=\operatorname{lcm}(\sigma)$.
  3. (3)
    The monomial $m$ is called a true gap of $\sigma$ if it is a gap of $\sigma$, and any bridge $m^{\prime}$ of $\sigma\cup m$ where $m\succ_{I}m^{\prime}$ is also a bridge of $\sigma$.
In [9, Theorem 2.24], the first two authors characterized type-1, potentially-type-2, and type-2 cells utilizing the concepts of bridges and true gaps. We focus primarily on the characterization of potentially-type-2 cells. Therefore, we present a slightly adjusted version of [9, Theorem 2.24 (b)] in the following.
Remark 2.9.
In a Barile-Macchia matching of $I$, a cell $\sigma$ is potentially-type-2 if and only if $m\succ_{I}\operatorname{sb}(\sigma)$ for each true gap $m$ of $\sigma$.
Based on the characterizations of type-1 and potentially-type-2 cells in [9, Theorem 2.24], we provide the following characterization of absolutely critical cells:
Corollary 2.10.
A cell is absolutely critical if and only if it has neither bridges nor true gaps.
In [9], the first two authors introduced a pivotal class of ideals called ”bridge-friendly” for analyzing the minimality of Barile-Macchia resolutions. As established in [9, Theorem 2.29], bridge-friendliness is a sufficient condition for an ideal to have a minimal Barile-Macchia resolution. We revisit the definition of this class of ideals using the concept of absolutely critical cells.
Definition 2.11.
[9, Definition 2.27] A monomial ideal $I$ is called bridge-friendly if, for some total order $\succ_{I}$ on ${\mathcal{G}}(I)$, every potentially-type-2 cell is of type-2 with respect to ($\succ_{I}$). Equivalently, all $A$-critical cells of $I$ are absolutely critical. Here, $A$ represents the Barile-Macchia matching of $I$ with respect to ($\succ_{I}$).
In the next chapter, we study a class of ideals with the bridge-friendly property.
3. Minimal free resolutions of path ideals of paths
In this section, our primary goal is to investigate the bridge-friendliness and, consequently, the minimal Barile-Macchia resolutions of path ideals of paths.
[…]
$I_{p}(L_{N})=(x_{1}x_{2}\cdots x_{p},\ x_{2}x_{3}\cdots x_{p+1},\ \cdots,\ x_{n}x_{n+1}\cdots x_{n+p-1}).$
Remark 3.1.
Path ideals can be viewed as an extension of edge ideals of graphs. Specifically, the $2$-path ideal of a graph coincides with its edge ideal.
The set of minimal generators of the $p$-path ideal of $L$ is
${\mathcal{G}}(I_{p}(L_{N}))=\{m_{1},m_{2},\ldots,m_{n}\}$
where $m_{i}:=x_{i}x_{i+1}\cdots x_{i+p-1}$ for each $1\leq i\leq n$. Fix a total order ()succeeds($\succ$)( ≻ ) on ${\mathcal{G}}(I_{p}(L_{N}))$ such that
$m_{1}\succ m_{2}\succ\cdots\succ m_{n}.$
Throughout the rest of this chapter, our focus is on the monomial ideal $I_{p}(L_{N})$. In particular, we examine its Barile-Macchia matching and resolution with respect to the aforementioned total order. For ease of readability, we introduce the following notation, which is consistently employed throughout the paper.
Notation 3.2.
For each $1\leq i\leq n$, let $M_{i}$ denote the set of all monomials in ${\mathcal{G}}(I_{p}(L_{N}))$ that are divisible by $x_{i}$. In other words,
$M_{i}=\{m_{i-p+1},\ldots,m_{i-1},m_{i}\}.$
Note that there are less than $p$ monomials in $M_{i}$ when $i\leq p-1$.
We begin our analysis with the following lemma, which serves as a foundational tool for the classification of bridges, gaps, and true gaps of a subset $\sigma$ of ${\mathcal{G}}(I_{p}(L_{N}))$.
Lemma 3.3.
Let $\sigma$ be a cell of $I_{p}(L_{N})$ and consider a monomial $m_{i}\in{\mathcal{G}}(I_{p}(L_{N}))$ such that $m_{i}\notin\sigma$. Assume that $\operatorname{lcm}(\sigma)$ is divisible by $m_{i}$. Then there exist monomials $m_{j}\in\sigma\cap M_{i}$ and $m_{k}\in\sigma\cap M_{i+p-1}$ such the distance between these monomials along $m_{i}$ is at most $p$, i.e., $k-j\leq p$.
Proof.
Since $\operatorname{lcm}(\sigma)$ is divisible by $m_{i}$, there exist monomials $m_{j}\in\sigma\cap M_{i}$ and $m_{k}\in\sigma\cap M_{i+p-1}$. Pick the largest such $j$ and smallest such $k$. Then any factor of $m_{i}$ divides either $m_{j}$ or $m_{k}$. In other words, $m_{i}$ divides $\operatorname{lcm}(m_{j},m_{k})$. For the sake of contradiction, assume that $k-j>p$. Consider the variable $x_{j+p}$. Note that $x_{j+p}$ does not divide neither $m_{j}$ nor $m_{k}$, where the latter comes from the assumption that $j+p<k$. On the other hand, since $m_{j}\in M_{i}$ and $m_{k}\in M_{i+p-1}$, we have $i+1\leq j+p<k\leq i+p-1$, which implies that $x_{j+p}$ divides $m_{i}$, a contradiction. Thus, we conclude that $k-j\leq p$. ∎
From Lemma 3.3, we can directly derive the characterization of bridges and gaps. It is important to note that in any cell $\sigma$ of $I_{p}(L_{N})$, neither $m_{1}$ nor $m_{n}$ can be a gap or a bridge.
Proposition 3.4.
Let $\sigma$ be a cell of $I_{p}(L_{N})$. For a monomial $m_{i}\in{\mathcal{G}}(I_{p}(L_{N}))$ with $1<i<n$, the following statements hold:
  1. (1)
    $m_{i}$ is a bridge of $\sigma$ if and only if $m_{i}\in\sigma$ and there exist monomials $m_{j}\in\sigma\cap M_{i}$ and $m_{k}\in\sigma\cap M_{i+p-1}$ such that $j<i<k$ and $k-j\leq p$.
  2. (2)
    $m_{i}$ is a gap of $\sigma$ if and only if $m_{i}\notin\sigma$ and there exist monomials $m_{j}\in\sigma\cap M_{i}$ and $m_{k}\in\sigma\cap M_{i+p-1}$ such that $k-j\leq p$. In particular, we have $j<i<k$.
Remark 3.5.
The monomials $m_{j}$ and $m_{k}$, which are highlighted in the proof of Lemma 3.3 and mentioned in the statement of Proposition 3.4, can be chosen to be those closest to $m_{i}$ among the monomials in $\sigma\setminus m_{i}$ in terms of their distance to $m_{i}$.
We present a characterization of true gaps below, which proves to be more intricate than the characterization of bridges and gaps.
Proposition 3.6.
Let $\sigma$ be a cell of $I_{p}(L_{N})$ and consider a monomial $m_{i}\in{\mathcal{G}}(I_{p}(L_{N}))$ where $1<i<n$. Assume $m_{i}$ does not dominate any bridges of $\sigma$. Then the monomial $m_{i}$ is a true gap of $\sigma$ if and only if the following statements hold:
  1. (a)
    $m_{i}\notin\sigma$ and there exist monomials $m_{j}\in\sigma\cap M_{i}$ and $m_{k}\in\sigma\cap M_{i+p-1}$ such that $k-j\leq p$.
  2. (b)
    $\sigma\cap M_{i+p-1}=\{m_{k}\}$ where $m_{k}$ is the element from part (a).
  3. (c)
    If $i+p\leq n$, then $m_{i+p}\notin\sigma$.
Proof.
For the forward direction, suppose $m_{i}$ is a true gap of $\sigma$. In order to prove (a), (b), and (c), it is useful to recall that $m_{i}$ is a true gap $\sigma$ that does not dominate any bridges of $\sigma$ if and only if $m_{i}$ is a gap of $\sigma$ and $\operatorname{sb}(\sigma\cup m_{i})=m_{i}$ by [9, Proposition 2.21].
  1. (a)
    By the definition of a true gap, $m_{i}$ is a gap. Hence, (a) holds by Proposition 3.4 (2).
  2. (b)
    It follows from (a) that $m_{k}\in\sigma\cap M_{i+p-1}$ for $i<k\leq n$. Suppose, for the sake of contradiction, that there exists another monomial $m_{t}$ in $\sigma\cap M_{i+p-1}$. Referring to Remark 3.5, without loss of generality, we may assume $k$ is the smallest index of a monomial in $\sigma\cap M_{i+p-1}$. So, $k<t$. Observe that $m_{k}$ divides $\operatorname{lcm}(m_{i},m_{t})$, implying that $m_{k}$ is a bridge of $\sigma\cup m_{i}$. This contradicts the condition $\operatorname{sb}(\sigma\cup m_{i})=m_{i}$. Thus, $m_{k}$ is the only monomial in $\sigma\cap M_{i+p-1}$.
  3. (c)
    Assume $i+p\leq n$. Since $\operatorname{sb}(\sigma\cup m_{i})=m_{i}$, the monomial $m_{k}$ is not a bridge of $\sigma\cup m_{i}$. In particular, this implies that $m_{i+p}\notin\sigma$.
For the other direction, assume the conditions (a), (b), and (c) hold. For contradiction, suppose that $m_{i}$ is not a true gap of $\sigma$. Since $m_{i}$ is a gap of $\sigma$ by (a), $m_{i}$ is not a true gap of $\sigma$ if and only if $m_{i}\succ m_{t}\coloneqq\operatorname{sb}(\sigma\cup m_{i})$ by [9, Proposition 2.21]. In particular, this implies that $i<t$. Since $m_{i}$ does not dominate any bridges of $\sigma$, $m_{t}$ is not a bridge of $\sigma$. According to Proposition 3.4 (1), we have $m_{t}\in\sigma$, and there are monomials $m_{j_{t}},m_{k_{t}}\in\sigma\cup m_{i}$ such that $j_{t}<t<k_{t}$ and $k_{t}-j_{t}\leq p$. Since $m_{t}$ is not a bridge of $\sigma$, either $j_{t}$ or $k_{t}$ has to be $i$. Given that $i<t$, the feasible scenario is $j_{t}=i$. This situation means that $m_{t}$ is an element of $\sigma\cap M_{i+p-1}$ as $i<t<k_{t}\leq i+p$. The condition (b) implies that $m_{k_{t}}$ is in $\sigma$ but is not contained in $M_{i+p-1}$. Hence, $k_{t}=i+p$, which leads to $m_{k_{t}}=m_{i+p}\in\sigma$, a direct contradiction to condition (c). As a result, $m_{i}$ must indeed be a true gap of $\sigma$. ∎
Example 3.7.
Consider the $3$-path ideal of an $8$-path,
$I=(x_{1}x_{2}x_{3},x_{2}x_{3}x_{4},x_{3}x_{4}x_{5},x_{4}x_{5}x_{6},x_{5}x_{6}x_{7},x_{6}x_{7}x_{8}).$
Consider the subset $\sigma=\{m_{1},m_{4},m_{6}\}$. It has no bridges, and its gaps are $m_{2},m_{3}$, and $m_{5}$. Using Proposition 3.6, we identify which among these gaps are true gaps.
  • The monomial $m_{2}$ is a true gap of $\sigma$. This is confirmed by: (a) the observation that $m_{2}\notin\sigma$ and $m_{1},m_{4}\in\sigma$, with the difference in their indices satisfying $4-1\leq 3$; (b) within the set $M_{4}=\{m_{2},m_{3},m_{4}\}$, only $m_{4}$ belongs to $\sigma$; and (c) the monomial $m_{2+3}=m_{5}$ is not in $\sigma$.
  • The monomial $m_{3}$ is not a true gap of $\sigma$. This is due to the failure of part (c) of Proposition 3.6 (3), given that $m_{6}\in\sigma$.
  • The monomial $m_{5}$ is a true gap of $\sigma$ based on: (a) $m_{5}\notin\sigma$ and both $m_{4}$ and $m_{6}$ are in $\sigma$, satisfying $6-4\leq 3$; (b) from the set $M_{7}=\{m_{5},m_{6}\}$, only $m_{6}$ is in $\sigma$. Furthermore, (c) is not applicable since $8>6$, and thus, $m_{5+3}=m_{8}$ does not exist.
In what follows, we show that path ideals of paths are bridge-friendly.
Theorem 3.8.
The path ideal $I_{p}(L_{N})$ is bridge-friendly, and its Barile-Macchia resolution is  minimal.
Proof.
To demonstrate the bridge-friendliness of $I_{p}(L_{N})$, we use [9, Lemma 2.33]. Specifically, our goal is to establish that, for any potentially-type-2 cell $\sigma$ (should it exist), there does not exist an $m\in{\mathcal{G}}(I_{p}(L_{N}))$ such that $m$ is a true gap of $\sigma\setminus\operatorname{sb}(\sigma)$ and that $\operatorname{sb}(\sigma)\succ m$.
If $I_{p}(L_{N})$ lacks potentially-type-2 cells, then its Taylor resolution is inherently minimal, making the path ideal bridge-friendly. So, we may assume $\sigma$ has potentially-type-2 cells. Now, consider a potentially-type-2 cell $\sigma$. This means $\operatorname{sb}(\sigma)$ exists. Note that $\operatorname{sb}(\sigma)\neq m_{n}$ since $m_{n}$ cannot be a bridge of $\sigma$. Then, there exists a monomial $m_{i}\in{\mathcal{G}}(I_{p}(L_{N}))$ satisfying $\operatorname{sb}(\sigma)\succ m_{i}$. We claim that $m_{i}$ is not a true gap of $\sigma\setminus\operatorname{sb}(\sigma)$. Suppose, for the sake of contradiction, that $m_{i}$ is a true gap. We first note that $m_{i}$ does not dominate any bridges of $\sigma$ due to the fact that $\operatorname{sb}(\sigma)\succ m_{i}$. Hence, by Proposition 3.6, we deduce that $m_{i}$ is a true gap of $\sigma$. The detailed justification is omitted as it directly arises from the application of Proposition 3.6 to our current hypothesis - that $m_{i}$ is a true gap of $\sigma\setminus\operatorname{sb}(\sigma)$, combined with our other hypothesis that $\operatorname{sb}(\sigma)\succ m_{i}$. This scenario implies that every bridge of $\sigma$ dominates a true gap, contradicting Remark 2.9 since $\sigma$ is potentially-type-2. Thus, no such $m_{i}$ exists, and by [9, Lemma 2.33], the path ideal $I_{p}(L_{N})$ is bridge-friendly.
The minimality of the Barile-Macchia resolution is a direct consequence of [9, Theorem 2.26]. ∎
In the subsequent discussion, our goal is to demonstrate that for every multidegree $m$, i.e., a monomial in $R$, there exists at most one critical cell $\sigma$ such that $\operatorname{lcm}(\sigma)=m$. Establishing the uniqueness of this critical cell for each multidegree allows us to compute both the projective dimension and Betti numbers of the path ideal via its Barile-Macchia resolution. To pave the way for this claim, we first introduce several auxiliary lemmas addressing the necessary technical details.
Before delving into these lemmas, it is worth recalling [9, Corollary 2.28] which characterizes the critical cells of $I_{p}(L_{N})$ by their lack of bridges and true gaps, a consequence of $I_{p}(L_{N})$ being bridge-friendly.
Lemma 3.9.
Let $\sigma$ be a critical cell of $I_{p}(L_{N})$, and let $x_{i}$ be a variable that divides $\operatorname{lcm}(\sigma)$. If none of the monomials $m_{k}$ with $k\in\{i-p+2,\ldots,i\}$ are in $\sigma$, then $m_{i-p+1}\in\sigma$.
Proof.
Note that $\sigma\cap M_{i}\neq\emptyset$ since $x_{i}$ divides $\operatorname{lcm}(\sigma)$. Given the conditions of the lemma, it is immediate that $m_{i-p+1}\in\sigma$. ∎
Lemma 3.10.
Let $\sigma$ be a critical cell of $I_{p}(L_{N})$, and $b\in\mathbb{N}\cup\{\infty\}$. Assume that there exists an integer $b^{\prime}\leq b-p-1$ such that among the set of monomials $\{m_{b^{\prime}-p+1},m_{b^{\prime}-p+2},\ldots,m_{b-p-1}\}$, only $m_{b^{\prime}-p+1}$ belongs to $\sigma$. Then the following statements hold:
  1. (1)
    No monomial in the set $\{m_{b^{\prime}-2p+1},\ldots,m_{b^{\prime}-p-1}\}$ belongs to $\sigma$.
  2. (2)
    $\operatorname{lcm}(\sigma)$ is divisible by $x_{b^{\prime}-p}$ if and only if $m_{b^{\prime}-p}$ belongs to $\sigma$.
Proof.
Suppose, for the sake of contradiction, that (1) does not hold, i.e., there exists a monomial $m_{k}$ in $\sigma\cap\{m_{b^{\prime}-2p+1},\ldots,m_{b^{\prime}-p-1}\}$. Under this assumption, we show that $m_{b^{\prime}-p}$ is either a bridge or a true gap of $\sigma$ which leads to a contradiction because $I_{p}(L_{N})$ is bridge-friendly and $\sigma$ has no bridges or true gaps by [9, Corollary 2.28].
Observe that $m_{b^{\prime}-p}$ divides $\operatorname{lcm}(\sigma\setminus m_{b^{\prime}-p})$. This is because both $m_{k}$ and $m_{b^{\prime}-p+1}$ are in $\sigma$ and the difference $(b^{\prime}-p+1)-k$ is at most $p$. Hence, $m_{b^{\prime}-p}$ is either a bridge or a gap of $\sigma$ by Proposition 3.4. If $m_{b^{\prime}-p}$ is a bridge of $\sigma$, we are done as discussed above. So, assume it is a gap of $\sigma$. Note that $m_{b^{\prime}-p}$ does not dominate any bridges of $\sigma$. Under these assumptions, $m_{b^{\prime}-p}$ is a true gap of $\sigma$ by Proposition 3.6 as its conditions (a), (b) and (c) are met:
  1. (a)
    $m_{b^{\prime}-p}$ is a gap of $\sigma$.
  2. (b)
    It follows from the hypothesis of the lemma that the only monomial in $\sigma\cap M_{b^{\prime}}$ is $m_{b^{\prime}-p+1}$.
  3. (c)
    Following the previous point, it is clear that $m_{b^{\prime}}$ does not belong to $\sigma$ (assuming it exists).
This completes the proof of (1). In order to prove (2), observe that $m_{b^{\prime}-p}\in\sigma$ implies $x_{b^{\prime}-p}$ divides $\operatorname{lcm}(\sigma)$. On the other hand, if $\operatorname{lcm}(\sigma)$ is divisible by $x_{b^{\prime}-p}$, then $\sigma\cap M_{b^{\prime}-p}$ is nonempty. By (1), the only such monomial in $\sigma\cap M_{b^{\prime}-p}$ is $m_{b^{\prime}-p}$. ∎
Lemma 3.11.
Let $\sigma$ be a critical cell of $I_{p}(L_{N})$ and $b\in\mathbb{N}\cup\{\infty\}$. Define the number $b^{\prime}$ as:
$b^{\prime}\coloneqq\begin{cases}\sup\{i\colon x_{i}\mid\operatorname{lcm}(\sigma)\text{ and }i\leq b-p-1\}&\text{if such }i\text{ exists},\\ -\infty&\text{otherwise}.\end{cases}$
Assume none of the monomials in the set $\{m_{b-2p+1},\ldots,m_{b-p-1}\}$ belong to $\sigma$ when $b$ is finite. Then among the monomials from the set $\{m_{b^{\prime}-p+1},m_{b^{\prime}-p+2},\ldots,m_{b-p-1}\}$, only $m_{b^{\prime}-p+1}$ belongs to $\sigma$.
Proof.
We analyze the two possible cases for $b$ separately.
Case 1: $b=\infty$. By the definition of $b^{\prime}$, it is clear that $b^{\prime}$ is finite. Given this, for any integer $k\geq b^{\prime}+1$, $\operatorname{lcm}(\sigma)$ is not divisible by $x_{k}$. This immediately implies that $\sigma\cap\{m_{b^{\prime}-p+2},\ldots,m_{k}\}$ is empty. Moreover, as $x_{b^{\prime}}$ divides $\operatorname{lcm}(\sigma)$, Lemma 3.9 guarantees that $m_{b^{\prime}-p+1}$ belongs to $\sigma$.
Case 2: $b\in\mathbb{N}$. From the definition of $b^{\prime}$, it is straightforward to see that $b^{\prime}\leq b-p-1$. If $b^{\prime}<b-p-1$, it can be deduced that $\operatorname{lcm}(\sigma)$ is not divisible by any variable from the set $\{x_{b^{\prime}+1},\ldots,x_{b-p-1}\}$. Consequently, $\sigma\cap\{m_{b^{\prime}-p+2},\ldots,m_{b-p-1}\}$ is empty. With the divisibility of $\operatorname{lcm}(\sigma)$ by $x_{b^{\prime}}$, Lemma 3.9 indicates that $m_{b^{\prime}-p+1}$ is indeed in $\sigma$. On the other hand, if $b^{\prime}=b-p-1$, then by the definition of $b^{\prime}$, we know that $x_{b-p-1}$ divides $\operatorname{lcm}(\sigma)$. Since $\sigma\cap\{m_{b-2p+1},\ldots,m_{b-p-1}\}$ is empty by the assumption of this lemma, the monomial $m_{b^{\prime}-p+1}=m_{b-2p}$ belongs to $\sigma$ by Lemma 3.9. ∎
Building upon the preceding three lemmas, we introduce a sequence that is key to understanding the content of a given critical subset $\sigma$. Here, by “content of $\sigma$”, we refer to the collection of monomials from ${\mathcal{G}}(I_{p}(L_{N}))$ that are in $\sigma$.
Proposition 3.12.
For a critical cell $\sigma$ of $I_{p}(L_{N})$, define the sequence $\{b_{i}\}_{i\in\mathbb{N}\cup\{0\}}$ by setting $b_{0}=\infty$ and, for $i\geq 1$,
$b_{i}=\sup\{j:x_{j}\mid\operatorname{lcm}(\sigma)\text{ and }j\leq b_{i-1}-p-1\},$
with $b_{i}=-\infty$ if no such $j$ exists. Then:
  1. (1)
    There exists $l\geq 1$ such that $b_{l}\neq-\infty$ and $b_{l+1}=-\infty$.
  2. (2)
    When $1\leq i\leq l$, only $m_{b_{i}-p+1}$ from the set $\{m_{b_{i}-p+1},\ldots,m_{b_{i-1}-p-1}\}$ is in $\sigma$.
  3. (3)
    When $i=l+1$, none of the monomials $m_{k}$ with $k\leq b_{l}-p-1$ are in $\sigma$.
Proof.
Given that $\sigma$ is a critical cell of $I_{p}(L_{N})$, it is non-empty, ensuring $b_{1}$ is finite. As the sequence $\{b_{i}\}$ decreases strictly after $b_{1}$ until it reaches $\infty$- ∞, there must be an $l\geq 1$ for which $\{i:x_{i}\mid\operatorname{lcm}(\sigma)\text{ and }i\leq b_{l-1}-p-1\}\neq\emptyset$ but $\{i:x_{i}\mid\operatorname{lcm}(\sigma)\text{ and }i\leq b_{l}-p-1\}=\emptyset$. This confirms that $b_{l}\neq-\infty$ and $b_{l+1}=-\infty$.
For $1\leq i\leq l$, we use induction on $i$. Lemma 3.11 covers our base case. For a fixed $i<l$, suppose $\sigma\cap\{m_{b_{j}-p+2},\ldots,m_{b_{j-1}-p-1}\}$ is empty and $m_{b_{j}-p+1}\in\sigma$ for each $j\leq i$. The induction hypothesis implies that only $m_{b_{i}-p+1}$ from the set $\{m_{b_{i}-p+1},\ldots,m_{b_{i-1}-p-1}\}$ belongs to $\sigma$. By Lemma 3.10, $\sigma\cap\{m_{b_{i}-2p+1},\ldots,m_{b_{i}-p-1}\}$ is empty. Given the constraints on $b_{i}$ and $b_{i+1}$, Lemma 3.11 implies that only $m_{b_{i+1}-p+1}$ from the set $\{m_{b_{i+1}-p+1},\ldots,m_{b_{i}-p-1}\}$ is in $\sigma$.
For the case $i=l+1$, note that no $x_{i}$ divides $\operatorname{lcm}(\sigma)$ for $i\leq b_{l}-p-1$. Thus, $\operatorname{lcm}(\sigma)$ is not divisible by any variable $x_{k}$ with $k\leq b_{l}-p-1$. This implies that $\sigma\cap\{m_{k-p+1},\ldots,m_{k}\}$ is empty for any $k$. Hence, $m_{k}$ is not in $\sigma$ for any $k\leq b_{l}-p-1$. ∎
Proposition 3.13.
For two critical cells $\sigma$ and $\sigma^{\prime}$ of $I_{p}(L_{N})$, if $\operatorname{lcm}(\sigma)=\operatorname{lcm}(\sigma^{\prime})$, then $\sigma=\sigma^{\prime}$.
Proof.
Assume that $\operatorname{lcm}(\sigma)=\operatorname{lcm}(\sigma^{\prime})$. Then, the decreasing sequences defined for $\sigma$ and $\sigma^{\prime}$ in Proposition 3.12 are identical. This is because the values of $b_{i}$ of a cell are determined solely by the least common multiple of the cell.
Based on Proposition 3.12, the monomial $m_{{b_{i}}-p+1}$ is present in both $\sigma$ and $\sigma^{\prime}$ for all $1\leq i\leq l$. The only remaining monomials in ${\mathcal{G}}(I_{p}(L_{N}))$ that could potentially belong to either $\sigma$ or $\sigma^{\prime}$ are those denoted by $m_{{b_{1}}-p},\ldots,m_{{b_{l}}-p}$. We can justify this claim with both Proposition 3.12 and the following decomposition of the real line:
$\mathbb{R}=$ $\,(b_{l+1},b_{l}]\cup\left(\bigcup_{i=1}^{l-1}(b_{i+1},b_{i}]\right)\cup(b_{1},b_{0})$
$=$ $\,(b_{l+1},b_{l}-p+1]\cup\left(\bigcup_{i=1}^{l-1}(b_{i+1}-p+1,b_{i}-p+1]\right)\cup(b_{1}-p+1,b_{0}).$
Considering Lemma 3.10, which states that $m_{b_{i}-p}$ is contained in the critical cell $\tau$ if and only if $x_{b_{i}-p}$ divides $\operatorname{lcm}(\tau)$ for $\tau\in\{\sigma,\sigma^{\prime}\}$, we conclude that $\sigma=\sigma^{\prime}$. ∎
From the preceding proposition, we deduce that the least common multiples of distinct critical cells of $I_{p}(L_{N})$ are different. This particularly implies the following information on its multi-graded Betti numbers.
Corollary 3.14.
For any monomial $m$ and any integer $i$, we have
$\beta_{i,m}(R/I_{p}(L_{N}))=\begin{cases}1&\text{ if there exists a critical subset of cardinality }i\text{ whose lcm equals }m,\\ 0&\text{ otherwise.}\end{cases}$
From the characterization of critical cells in Proposition 3.12 and the insights from the proof of Proposition 3.13, we can deduce the projective dimension of $I_{p}(L_{n+p-1})$. Specifically, among every collection of $p+1$ consecutive monomials, at most two can be in a critical cell. Consequently, we can derive the maximal cardinality of a critical cell, thus obtaining the projective dimension.
Corollary 3.15.
Let $n$ be expressed as $n=(p+1)q+s$ where $0\leq s\leq p$. The projective dimension of $R/I_{p}(L_{n+p-1})$ is given by:
$\operatorname{pdim}(R/I_{p}(L_{n+p-1}))=\begin{cases}2q&\text{if }s=0,\\ 2q+1&\text{if }s=1,\\ 2q+2&\text{otherwise}.\end{cases}$
A formula for the projective dimension of the path ideal of a path was also given in [2]. The formula from [2] matches ours. However, while we express the projective dimension based on the number of minimal generators of the $p$-path ideal, [2] does so using the length of the path.
We also recover the recursive formula for graded Betti numbers from [7]. This formula was utilized in [2] to provide explicit calculations for the projective dimension and regularity of path ideals of paths and cycles.
Theorem 3.16.
For all indices $r,d$, we have
$\beta_{r,d}(R/I_{p}(L_{n}))=\beta_{r,d}(R/I_{p}(L_{n-1}))+\beta_{r-1,d-{p}}(R/I_{p}(L_{n-(p+1)}))+\beta_{r-2,d-(p+1)}(R/I_{p}(L_{n-(p+1)})).$
Proof.
By Theorem 3.14, $\beta_{r,d}(R/I_{p}(L_{n}))$ counts the critical cells of cardinality $r$ and degree $d$. To derive our desired expression, it suffices to partition the set of critical subsets of cardinality $r$ and degree $d$ in an appropriate way. Consider a critical cell $\sigma$ of $I_{p}(L_{n})$. The following three scenarios for $\sigma$ completes the proof:
  1. (1)
    If $m_{n-p+1}\notin\sigma$, then $\sigma$ is a critical cell of $I_{p}(L_{n})$ if and only if it is a critical cell of $I_{p}(L_{n-1})$.
  2. (2)
    If $m_{n-p+1}\in\sigma$ but $m_{n-p}\notin\sigma$, then $\sigma$ is a critical cell of $I_{p}(L_{n})$ if and only if $\sigma\backslash\{m_{n-p+1}\}$ is a critical cell of $I_{p}(L_{n-(p+1)})$ by Corollary 3.4 and Proposition 3.6.
  3. (3)
    If both $m_{n-p+1}$ and $m_{n-p}$ are in $\sigma$, then $\sigma$ is a critical cell of $I_{p}(L_{n})$ if and only if $\sigma\backslash\{m_{n-p+1},m_{n-p}\}$ is a critical cell of $I_{p}(L_{n-(p+1)})$ by Corollary 3.4 and Proposition 3.6. ∎
Barile-Macchia resolutions are cellular, i.e., they are supported on CW complexes. We first remark that the dimension of the CW complex that supports the minimal resolution of a monomial ideal equals its projective dimension. In general, the minimal resolution of any monomial ideal of projective dimension $1$ is supported on a tree [15, Theorem 1]. In fact, in the cases where the path ideals of paths have projective dimension $1$, their minimal resolutions are supported on paths, which can be shown using the techniques that will be employed in the next example. In what follows, we provide an example where the path ideal of a path has projective dimension 2.
Example 3.17.
Consider the path ideal $I=I_{p}(L_{2p+1})$ under the total order
$m_{1}\succ m_{2}\succ\cdots\succ m_{p+2}.$
For any subset $\sigma=\{m_{i_{1}},\ldots,m_{i_{k}}\}$, each element of $\sigma$ (except $m_{i_{1}}$ and $m_{i_{k}}$) is a bridge of $\sigma$ due to Proposition 3.4 (1). Consequently, $m_{i_{k-1}}$ emerges as the smallest bridge of $\sigma$. Therefore, the Barile-Macchia matching of $I$ with respect to ()succeeds($\succ$)( ≻ ), denoted by $A$, is achieved by removing the penultimate element at every iteration. Note that there is only one cell (of cardinality of at least 3) with no bridges: $\{m_{1},m_{p+1},m_{p+2}\}$. This results in the following list of all critical cells of $I$:
,$\emptyset,\{m_{1}\},\dots,\{m_{p+2}\},\{m_{1},m_{2}\},\{m_{2},m_{3}\},\dots,\{m_{p+1},m_{p+2}\},\{m_{p+2},m_{1}\},\{m_{1},m_{p+1},m_{p+2}\}.$
Given two distinct critical cells, $\sigma$ and $\sigma^{\prime}$, their least common multiples are different by Proposition 3.13. This uniqueness ensures that the Barile-Macchia resolution of $I$ is minimal by Theorem 2.2.
By discrete Morse theory, there exists a cellular complex that supports this minimal free resolution, and it has $(p+2)$ many 00-cells, $(p+2)$ many $1$-cells, and one $2$-cell. One obvious object that fits this description is the following $(p+2)$-gon:
$\{m_{1},m_{2}\}$$\{m_{2},m_{3}\}$$\{m_{3},m_{4}\}$$\{m_{4},m_{5}\}$$\{m_{1},m_{p+2}\}$$\{m_{1}\}$$\{m_{2}\}$$\{m_{3}\}$$\{m_{4}\}$$\{m_{5}\}$$\{m_{p+2}\}$$\{m_{1},m_{p+1}m,m_{p+2}\}$
One can show that this $(p+2)$-gon, denoted by $\Delta$, supports a free resolution of $R/I$. By [6, Lemma 2.2], it suffices to demonstrate that for each monomial $m$, the subcomplex $\Delta[m]\coloneqq\{\sigma\in\Delta\colon\operatorname{lcm}(\sigma)\text{ divides }m\}$ of $\Delta$ is either empty or contractable. By exhausting all possible cases, one can confirm that the subcomplex $\Delta[m]$ is either empty, a line segment, or the $(p+2)$-gon $\Delta$ itself for each squarefree monomial $m$. Thus, it is well-known that the latter two objects are both contractable.
4. Minimal free resolutions of path ideals of cycles
In this section, we turn our attention to path ideals of cycles, demonstrating that these ideals have minimal cellular resolutions. While in the preceding section we derived this outcome for path ideals of paths through identifying a minimal Barile-Macchia resolution of $R/I$ with respect to a specific total order on ${\mathcal{G}}(I)$, this approach falls short for path ideals of cycles. For instance, when we consider the edge ideal of a 9-cycle, it has no minimal Barile-Macchia resolutions as pointed out in [9, Remark 4.23].
In our investigation of path ideals of cycles, we transition our focus towards the generalized Barile-Macchia resolutions. These are Morse resolutions and can be considered as an extension of the Barile-Macchia resolutions, introduced in [9]. The crux of these resolutions lies in utilizing a collection of total orders on ${\mathcal{G}}(I)$, instead of one. For the reader’s convenience, we restate the construction of generalized Barile-Macchia resolutions from [9] along with a theorem stating that they induce cellular free resolutions. First, recall that $G_{X}$ is the directed graph obtained from the Taylor complex of $I$.
Theorem 4.1.
[9, Theorem 5.19] For a monomial $u\in R$, let $G_{u}$ be the induced subgraph of $G_{X}$ on the vertices $\sigma\subseteq{\mathcal{G}}(I)$ where $\operatorname{lcm}(\sigma)=u$. Consider a total order $(\succ_{u})$ on ${\mathcal{G}}(I)$ for each monomial $u\in R$.
Let $A$ be the union of all $A_{u}$, where $A_{u}$ is the collection of directed edges obtained by applying the Barile-Macchia Algorithm to $G_{u}$ with respect to $(\succ_{u})$ for each monomial $u\in R$. Then, $A$ is a homogeneous acyclic matching of $I$. The Morse resolution induced by $A$ is called the generalized Barile-Macchia resolution of $R/I$ with respect to $(\succ_{u})_{u\in R}$.
[…]
We begin our discussion with an observation on the construction of the set $A$, as outlined in Theorem 4.1, leveraging our previous findings on path ideals of paths.
Observation 4.2.
Let $A$ be a homogeneous acyclic matching constructed in accordance with Theorem 4.1 for the ideal $I_{p}(C_{n})$. To obtain a generalized Barile-Macchia resolution of $R/I_{p}(C_{n})$, it is necessary to determine its $A$-critical cells.
  1. (1)
    The following is immediate from the definition of $A$:
[…]
  2. (2)
    Assuming $u\neq x_{1}\cdots x_{n}$, remember that each vertex of $G_{X}$ is a subset of ${\mathcal{G}}(I_{p}(C_{n}))$. Note that $V(G_{u})$, the vertex set of $G_{u}$, is empty or every vertex of $G_{u}$ is a subset of ${\mathcal{G}}(J)$ where $J$ is the path ideal of some path. When $V(G_{u})$ is non-empty, Proposition 3.13 assures the existence of a total order ()succeeds($\succ$)( ≻ ) on ${\mathcal{G}}(J)$ that allows for precisely one $A$-critical cell.
In light of the above observation, our attention is on the $A_{u}$-critical cells where $u=x_{1}\cdots x_{n}$ to derive a minimal generalized Barile-Macchia resolution of $R/I_{p}(C_{n})$.
For the remainder of this section, let $u=x_{1}\cdots x_{n}$ and adopt the following total order ()succeeds($\succ$)( ≻ ) for  $u$:
$m_{1}\succ m_{2}\succ\cdots\succ m_{n}.$
Our primary objective is to demonstrate that all $A_{u}$-critical cells have the same cardinality. Consequently, the resulting generalized Barile-Macchia resolution is minimal by Theorem 2.2 and 4.2.
As a preliminary step, we examine the structure of $\sigma\in V(G_{u})$. Recall that $\operatorname{lcm}(\sigma)=u$ for each vertex $\sigma$ in $G_{u}$.
Proposition 4.3.
Let $\sigma\in V(G_{u})$. Then $\sigma$ is of the form $\{m_{i_{1}},\dots,m_{i_{t}}\}\subseteq{\mathcal{G}}(I_{p}(C_{n}))$ with the property that the distance between consecutive elements of $\sigma$ is at most $p$, i.e., $i_{j+1}-i_{j}\leq p$ for all $j\in\{1,\ldots,t-1\}$ and $i_{1}+n-i_{t}\leq p$.
Proof.
Suppose that either $i_{j+1}-i_{j}>p$ for some $j\in\{1,\ldots,t-1\}$ or $i_{1}+n-i_{t}>p$. In both cases, this implies that $u=x_{1}\cdots x_{n}$ cannot be divisible by $x_{i_{j}+p}$ or $x_{i_{t}+p}$, respectively, leading to a contradiction. ∎
Building upon our earlier observations, a key point we aim to elucidate is that every critical cell contains $m_{n}$.
Lemma 4.4.
If $\sigma$ is an $A_{u}$-critical cell, then $m_{n}$ belongs to $\sigma$.
Proof.
Suppose for contradiction that $\sigma$ is an $A_{u}$-critical cell that does not contain $m_{n}$. By Proposition 4.3, $m_{n}$ is a gap of $\sigma$. In particular, this implies that $m_{n}=\operatorname{sb}(\sigma\cup m_{n})$. Thus, $\sigma\cup m_{n}$ is potentially-type-2 by Remark 2.9. Moreover, it is not possible to find a cell $\tau$ such that $\tau\setminus\operatorname{sb}(\tau)=\sigma$ where $m_{n}\succ\operatorname{sb}(\tau)$. Hence, we can infer that the directed edge $(\sigma\cup m_{n},\sigma)$ belongs to $A_{u}$. This contradicts the assumption that $\sigma$ is an $A_{u}$-critical cell. Thus, $m_{n}\in\sigma$. ∎
Notation 4.5.
Similar to Notation 3.2 for paths, define $M_{i}$ to be the collection of all monomials in ${\mathcal{G}}(I_{p}(C_{n}))$ that are divisible by $x_{i}$. For any $i\geq p$, we have $M_{i}=\{m_{i-p+1},\ldots,m_{i}\}$. For any given $i$ such that $11\leq i\leq p-$ 1, we have
$M_{i}=\{m_{(n+i)-p+1},\ldots,m_{n},m_{1},\ldots,m_{i}\}.$
Clearly, $|M_{i}|=p$ for each value of $i$.
In the subsequent discussion, we analyze the vertices of $G_{u}$ and describe their bridges, gaps, and true gaps. Analogous to the path case, we can classify the bridges in a similar manner. The proof is omitted since it directly follows from Lemma 3.3, keeping in mind that indices are now considered modulo $n$. First, we consider gaps as their classification is immediate.
Proposition 4.6.
Let $\sigma\in V(G_{u})$. For $i\in\{1,\ldots,n\}$, a monomial $m_{i}$ is a gap of $\sigma$ if and only if $m_{i}\notin\sigma$.
Proof.
Note that adding $m_{i}$ to $\sigma$ does not change the least common multiple. Specifically, $\operatorname{lcm}(\sigma)=\operatorname{lcm}(\sigma\cup m_{i})=x_{1}\cdots x_{n}$, establishing the assertion. ∎
Proposition 4.7.
Let $\sigma\in V(G_{u})$, and $m_{i}\in{\mathcal{G}}(I_{p}(C_{n}))$ be a monomial. Then the monomial $m_{i}$ is a bridge of $\sigma$ if and only if the following conditions are met:
  1. (1)
    $m_{i}\in\sigma$.
  2. (2)
    There exist monomials $m_{j}$ and $m_{k}$ in $\sigma$ for $1\leq j<k\leq n$ with $i\neq j,k$ such that the distance between these two monomials along $m_{i}$ is at most $p$, i.e.,
    1. (a)
      $k-j\leq p$ when $j<i<k$.
    2. (b)
      $j+n-k\leq p$ otherwise.
Remark 4.8.
As in Remark 3.5, the monomials $m_{j}$ and $m_{k}$ from Proposition 4.7 are often chosen to be those closest to $m_{i}$ among the monomials in $\sigma$ in terms of their distance to $m_{i}$.
We now turn our attention to characterizing the true gaps for vertices in $G_{u}$ that are $A_{u}$-critical.
Proposition 4.9.
Let $\sigma\in V(G_{u})$ be an $A_{u}$-critical cell and let $m_{i}\in{\mathcal{G}}(I_{p}(C_{n}))$. Assume $m_{i}$ does not dominate any bridge of $\sigma$. Then $m_{i}$ is a true gap of $\sigma$ if and only if the following statements hold:
  1. (a)
    $m_{i}\notin\sigma$.
  2. (b)
    If there exists a monomial $m_{k}\in\sigma\cap M_{i+p-1}$, then none of the other monomials in $M_{i+p-1}$ belongs to $\sigma$. Furthermore, $m_{i+p}\notin\sigma$.
  3. (c)
    If $11\leq i\leq p-$ 1 and $\sigma\cap\{m_{1},\ldots,m_{i-1}\}$ is empty, then the only monomial in $M_{i}$ that belongs to $\sigma$ is $m_{n}$. Furthermore, $m_{i-p}\notin\sigma$.
Proof.
Suppose $m_{i}$ does not dominate any bridges of $\sigma$, i.e., $\operatorname{sb}(\sigma)\succ m_{i}$. We start our proof by observing $m_{n}\in\sigma$ by Proposition 4.4 since $\sigma$ is an $A_{u}$-critical cell.
We first deal with the forward direction. Assume that $m_{i}$ is a true gap of $\sigma$. Then, $m_{i}$ is also a gap of $\sigma$ and $m_{i}\notin\sigma$ by Proposition 4.6, which verifies condition (a). For condition (b), the arguments align closely with those presented in Proposition 3.6 for parts (b) and (c). Thus, the details are omitted to avoid redundancy.
Now, to verify condition (c), assume that $11\leq i\leq p-$ 1 and that $\sigma\cap\{m_{1},\ldots,m_{i-1}\}$ is empty. Given that $m_{n}\in M_{i}$ for the stated range of $i$, there exists a monomial $m_{t}\in\sigma$ that also belongs to $M_{i}$, but with $t\notin\{1,\ldots,i\}$. In this scenario, unless $m_{t}=m_{n}$, the monomial $m_{n}$ would emerge as a bridge for $\sigma\cup m_{i}$. This implies that $\operatorname{sb}(\sigma\cup m_{i})=m_{n}$, a contradiction to [9, Proposition 2.21]. Hence, the only monomial in $\sigma\cap M_{i}$ is $m_{n}$. For the last segment of (c), if we assume $m_{i-p}\in\sigma$, it leads to the conclusion that $\operatorname{sb}(\sigma\cup m_{i})=m_{n}$. This is, yet again, contradictory to [9, Proposition 2.21], thereby confirming that $m_{i-p}\notin\sigma$.
For the reverse direction, assume conditions (a), (b), and (c) hold. If we argue by contradiction and assume that $m_{i}$ is not a true gap of $\sigma$, then there exists a monomial $m_{t}$ that is a bridge of $\sigma\cup m_{i}$ but is not a bridge of $\sigma$, with $m_{i}\succ m_{t}$. Using Proposition 4.7, we deduce $m_{t}\in\sigma$. Furthermore, there exist monomials $m_{j_{t}}$ and $m_{k_{t}}$ in $\sigma\cup m_{i}$ with $1\leq j_{t}<k_{t}\leq n$ and $t\notin\{j_{t},k_{t}\}$ such that
  1. (1)
    $k_{t}-j_{t}\leq p$ when $j_{t}<t<k_{t}$ or
  2. (2)
    $j_{t}+n-k_{t}\leq p$ otherwise.
As discussed in 4.5, we may assume that $j_{t}$ and $k_{t}$ are the indices of monomials in $\sigma$ that are closest to $m_{t}$. Because $m_{t}$ is not a bridge of $\sigma$, either $j_{t}=i$ or $k_{t}=i$. If $j_{t}<t<k_{t}$, we must have $j_{t}=i$ since $m_{i}\succ m_{t}$. The reasoning for this case mirrors the corresponding portion of Proposition 3.6, eventually leading to a contradiction.
Note that we must have $j_{t}<k_{t}<t$ given $i<t$ when we consider other orderings of $j_{t},t$, and $k_{t}$ on the cycle. Based on (2), we have:
($\star$) $j_{t}<k_{t}<t<j_{t}+n\leq k_{t}+p\leq n+p.$
In the scenario where $k_{t}=i$, we have $m_{j_{t}}\in\sigma$. and considering ($\star$4), $m_{t}$ is in $M_{i+p-1}$. However, due to condition (b) and the fact that $m_{j_{t}+n}\in\sigma$, it must be that $m_{j_{t}+n}=m_{i+p}$, a contradiction to (b). For the final case, we can assume $j_{t}=i$ and with $j_{t}$ and $k_{t}$ being the closest monomials in $\sigma$ to $m_{t}$, the intersection $\sigma\cap\{m_{1},\ldots,m_{i-1}\}$ is empty. Given ($\star$4), we deduce that $i\leq p$. Since the only monomial in $\sigma\cap M_{i}$ is $m_{n}$ by (c) and $(i+n)-p\leq k_{t}\leq n$, we have $m_{t}=m_{n}$ and $m_{k_{t}}=m_{i-p}$. Consequently, $m_{k_{t}}=m_{i-p}\in\sigma$, a contradiction to condition (c). Thus, $m_{i}$ is indeed a true gap of $\sigma$.∎
Having addressed the true gaps of $A_{u}$-critical cells in the preceding proposition, our focus now shifts to the distinct classes of critical cells introduced in Definition 2.7: the absolutely critical and the fortunately critical cells. As noted in Corollary 2.10, the absolutely critical cells are uniquely characterized by the lack of both bridges and true gaps. On the other hand, the fortunately critical cells stand out. Their first element serves as their smallest bridge while still having no true gaps.
Lemma 4.10.
Let $\sigma\in V(G_{u})$ be an $A_{u}$-critical cell. If $\sigma$ is fortunately critical, then:
  1. (1)
    The smallest bridge of $\sigma$ satisfies $\operatorname{sb}(\sigma)\succ m_{p}$.
  2. (2)
    If $\sigma=\{m_{i_{1}},\ldots,m_{i_{t}}\}$ where $i_{1}<\cdots<i_{t}$, then $\operatorname{sb}(\sigma)=m_{i_{1}}.$
Furthermore, no monomial from the set $\{m_{p},m_{p+1},\ldots,m_{n}\}$ serves as a bridge or a true gap of $\sigma$ (irrespective of whether $\sigma$ is fortunately or absolutely critical).
Proof.
Let $\sigma$ be a vertex in $G_{u}$ that is fortunately critical. By definition, there exists another vertex $\sigma^{\prime}$ such that $\sigma\backslash\operatorname{sb}(\sigma)=\sigma^{\prime}\backslash\operatorname{sb}(\sigma^{\prime})$ and $\operatorname{sb}(\sigma)\succ\operatorname{sb}(\sigma^{\prime})$.
We start by noting that the monomial $\operatorname{sb}(\sigma^{\prime})$ is a true gap of $\sigma\backslash\operatorname{sb}(\sigma)$ by [9, Proposition 2.22]. Given that every true gap of $\sigma$ dominates $\operatorname{sb}(\sigma)$ by Remark 2.9, it follows that $\operatorname{sb}(\sigma^{\prime})$ cannot be a true gap of $\sigma$. This implies that there exists a monomial $m_{i}\in\sigma^{\prime}$ for which $\operatorname{sb}(\sigma\cup\operatorname{sb}(\sigma^{\prime}))=m_{i}$, with $\operatorname{sb}(\sigma^{\prime})\succ m_{i}$ and $m_{i}$ not being a bridge of $\sigma$. By Proposition 4.7, there exist monomials $m_{j}$ and $m_{k}$ in $\sigma\cup\operatorname{sb}(\sigma^{\prime})$ with $1\leq j<k\leq n$ and $i\notin\{j,k\}$ such that either:
  1. (a)
    $k-j\leq p$ and $j<i<k$ holds, or
  2. (b)
    $j+n\leq k+p$.
Inevitably, we have $\{\operatorname{sb}(\sigma),\operatorname{sb}(\sigma^{\prime})\}=\{m_{j},m_{k}\}$ since $m_{i}$ is neither a bridge of $\sigma$ nor $\sigma^{\prime}$. We dissect this further: If $j<i<k$, then $\operatorname{sb}(\sigma^{\prime})=m_{j}$ since $\operatorname{sb}(\sigma^{\prime})\succ m_{i}\succ m_{k}$. In this case, it is immediate that $m_{k}\in\sigma^{\prime}$. Since $m_{j},m_{i},m_{k}\in\sigma^{\prime}$ and $k-j\leq p$ by (a), the monomial $m_{i}$ is a bridge of $\sigma^{\prime}$ by Proposition 4.7, which is a contradiction. Thus, (b) holds. Moreover, since $\{m_{j},m_{k}\}=\{\operatorname{sb}(\sigma),\operatorname{sb}(\sigma^{\prime})\}$ and both $\operatorname{sb}(\sigma),\operatorname{sb}(\sigma^{\prime})$ dominate $m_{i}$, we must have $j<k<i$. Consequently, $m_{j}=\operatorname{sb}(\sigma)$ and $m_{k}=\operatorname{sb}(\sigma^{\prime})$. Since $m_{n}\in\sigma\cup\operatorname{sb}(\sigma^{\prime})$, we can conclude that $m_{n}$ is a bridge of $\sigma\cup\operatorname{sb}(\sigma^{\prime})$ by Proposition 4.7 and (b). This means $m_{i}=m_{n}$.
  1. (1)
    To prove $\operatorname{sb}(\sigma)\succ m_{p}$, we first note that $\operatorname{sb}(\sigma)=m_{j}\succeq m_{p}$ due to the inequality $j+n\leq k+p\leq n+p$ by (b). If $j=p$, (b) from above implies that we have $n\leq k$, which is not possible. So, $\operatorname{sb}(\sigma)\succ m_{p}$.
  2. (2)
    Let $\sigma=\{m_{i_{1}},\ldots,m_{i_{t}}\}$ where $i_{1}<\cdots<i_{t}$. Our goal is to demonstrate $\operatorname{sb}(\sigma)=m_{i_{1}}$. On the contrary, suppose there is a monomial $m_{s}$ in $\sigma$ such that $m_{s}\succ\operatorname{sb}(\sigma)=m_{j}$. Then, by Proposition 4.7, monomial $m_{n}$ is a bridge of $\sigma^{\prime}$ since $m_{k},m_{n},$ and $m_{s+n}$ are all members of $\sigma^{\prime}$ and the inequality $s+n-k<p$ holds true by (b) above. This posits a contradiction, as it means $\operatorname{sb}(\sigma^{\prime})=m_{k}\succ m_{n}$ while $m_{n}$ is a bridge of $\sigma^{\prime}$. Thus, $\operatorname{sb}(\sigma)=m_{i_{1}}$.
For the final part of the statement, consider a vertex $\sigma$ in $G_{u}$. If $\sigma$ is absolutely critical, it lacks both bridges and true gaps, by Corollary 2.10, thereby satisfying the given statement. When $\sigma$ is fortunately critical, the statement remains valid due to $\operatorname{sb}(\sigma)\succ m_{p}$, and the fact that each true gap of $\sigma$ dominates $\operatorname{sb}(\sigma)$ by Remark 2.9. ∎
Our primary objective is to comprehensively identify every element within an $A_{u}$-critical cell. We begin our identification with a series of observations and initiate the process by pinpointing specific values of $j$ for which $m_{j}\in\sigma$.
Lemma 4.11.
Let $\sigma=\{m_{i_{1}},\ldots,m_{i_{t}}\}$ be an $A_{u}$-critical cell. Then
  1. (a)
    $i_{t}=n$, and $m_{n}$ is not a bridge for $\sigma$.
  2. (b)
    $i_{t-1}=n-k$ for some $1\leq k\leq p$.
  3. (c)
    $i_{t-2}<n-p$.
  4. (d)
    $i_{1}=p-k+1$ where $k$ is number from the expression of $i_{t-1}$ in (a).
Proof.
Assume that $\sigma=\{m_{i_{1}},\ldots,m_{i_{t}}\}$ is an $A_{u}$-critical cell.
  1. (a)
    We begin by noting that $m_{i_{t}}=m_{n}$ as per Lemma 4.4. If $\sigma$ is absolutely critical, then, according to Corollary 2.10, it has no bridges. On the other hand, for a fortunately critical $\sigma$, Lemma 4.10 dictates that $\operatorname{sb}(\sigma)=m_{i_{1}}$. Given that $m_{i_{1}}\succ m_{n}$, it is evident that $m_{n}$ is not a bridge for $\sigma$.
  2. (b)
    Recall from Proposition 4.3 that $i_{t}-i_{t-1}\leq p$. Then, $i_{t-1}=n-k$ for some $1\leq k\leq p$.
  3. (c)
    For the sake of contradiction, suppose $m_{n-j}\in\sigma$ for $1\leq k<j\leq p$. Then, $m_{n-k}$ emerges as a bridge for $\sigma$. However, this contradicts the nature of $\sigma$, as it is either absolutely critical (hence having no bridges) or fortunately critical (where $\operatorname{sb}(\sigma)=m_{i_{1}}$).
  4. (d)
    Our initial step is to derive $p-k+1\leq i_{1}\leq p$. The upper bound $i_{1}\leq p$ is a direct consequence of Proposition 4.3. To establish the lower bound, if there exists an $m_{i}\in\sigma$ for $i\leq p-k$, then $m_{n}$ is a bridge of $\sigma$ by Proposition 4.7. This assertion, however, yields a contradiction by (a). Thus, $p-k+1\leq i_{1}\leq p$.
    If $k=1$, then it is immediate that $i_{1}=p$, satisfying the statement of (d). For $2\leq k\leq p$, it remains to show $i_{1}\leq p-k+1$. An significant insight here is that by Proposition 4.9 (b), $m_{n-k+1}$ cannot be a true gap of $\sigma$ since $M_{p-k}$ contains both $m_{n-k}$ and $m_{n}$ from $\sigma$. This means $\sigma\cup m_{n-k+1}$ has a bridge $m_{b}\in\sigma$ such that $m_{n-k+1}\succ m_{b}$. Since the only monomial in $\sigma$ that is dominated by $m_{n-k+1}$ is $m_{n}$, we have $m_{b}=m_{n}$. Using Proposition 4.7 for the bridge $m_{b}$, we conclude $i_{1}\leq p-k+1$. ∎
To identify the other elements of $\sigma$, we examine them in relation to the possible values of $k$. Here, $n-k$ is the penultimate element of $\sigma$ when $1\leq k\leq p$. The next two lemmas address the $k=1$ case and the $2\leq k\leq p$ case separately due to nuanced variations in their proofs. Together, these lemmas give a complete overview of all $A_{u}$-critical cells.
Lemma 4.12.
Let $\sigma=\{m_{i_{1}},\ldots,m_{i_{t}}\}$ be an $A_{u}$-critical cell with $i_{t-1}=n-1$. Then, the distance between consecutive elements of $\sigma$ alternates between $1$ and $p$. Specifically,
$\sigma=\{m_{p},\ldots,m_{n-(p+1)},m_{n-(p+1)},m_{n-1},m_{n}\}.$
Moreover, we have $p\equiv n$ or $n-1\pmod{p+1}$.
Proof.
The main idea of the proof revolves around retracing our steps from $i_{t-1}$, pinpointing the preceding indices of elements in $\sigma$ until we arrive at $i_{1}=p$. A recurring and instrumental point from Proposition 4.3 to note during the proof is: $i_{j}-i_{j-1}\leq p$ for $m_{i_{j-1}},m_{i_{j}}\in\sigma$.
Before we identify elements of $\sigma$, note that $\sigma$ is an absolutely critical cell by Lemma 4.10 because $i_{1}=p$. Recall from Lemma 4.11 that $i_{t-2}\leq n-1-p$. The first step of our analysis is immediate as $i_{t-1}-i_{t-2}=p$ by Lemma 4.11 (b) and (c). If $i_{t-2}=i_{1}$, we are done; otherwise, we move on to $i_{t-3}$.
Next, we show that $i_{t-2}-i_{t-3}=1$ which is equivalent to obtaining $m_{n-p-2}\in\sigma$. For the sake of contradiction, suppose $m_{n-p-2}\notin\sigma$. Since $\sigma$ is absolutely critical, $m_{n-p-2}$ cannot be a true gap of $\sigma$. This means $\sigma\cup m_{n-p-2}$ has a new bridge $m_{b}\in\sigma$ such that $m_{n-p-2}\succ m_{b}$. Then, by Proposition 4.7, there exists $m_{s}\in\sigma$ such that the distance between $m_{s}$ and $m_{n-p-2}$ along $m_{b}$ is at most $p$. Given that $m_{n-p-2}\succ m_{b}$, by Proposition 4.7, we have either
  1. (1)
    $s-(n-p-2)\leq p$ when $1\leq n-p-2<b<s\leq n$, or
  2. (2)
    $s+n-(n-p-2)\leq p$ when $1\leq s<n-p-2<b\leq n.$
The latter case is not possible because it implies that $s\leq-2$. Thus, $m_{b}=m_{n-1}$ and $m_{s}=m_{n}$. On the other hand, by (1), the distance between $m_{n-p-2}$ and $m_{n}$ must be at most $p$ but it is $p+2$, a contradiction. Thus, we conclude that $i_{t-3}=n-p-2$. If $i_{t-3}=i_{1}$, our task is complete; otherwise, our focus shifts to $i_{t-4}$.
Towards showing $i_{t-3}-i_{t-4}=p$, recall that $i_{t-3}-i_{t-4}\leq p$, or equivalently, $i_{t-4}\geq n-2(p+1)$. If $m_{n-2(p+1)}\notin\sigma$, then $i_{t-2}-i_{t-4}\leq p$, implying that $m_{i_{t-3}}$ is a bridge of $\sigma$ Proposition 4.7. This leads to a contradiction because $\sigma$ is absolutely critical. Hence, $i_{t-4}=n-2(p+1)$. If $i_{t-4}=i_{1}$, our investigation is complete; otherwise, we continue in the same fashion for $i_{t-5}$.
Subsequent distances between consecutive elements of $\sigma$ can be concluded by employing similar arguments in an alternating way until we reach $i_{1}=p$. This results in the congruence $p\equiv n\text{ or }n-1\pmod{p+1}$, concluding the proof. ∎
Lemma 4.13.
Let $\sigma=\{m_{i_{1}},\ldots,m_{i_{t}}\}$ be an $A_{u}$-critical cell with $i_{t-1}=n-k$ for some $2\leq k\leq p$. Then, the distance between consecutive elements of $\sigma$ alternates between $k$ and $(p+1)-k$. Specifically,
$\sigma=\{m_{p-k+1},\ldots,m_{n-k-(p+1)},m_{n-(p+1)},m_{n-k},m_{n}\}$
Moreover, we have $p-k+1\equiv n$ or $n-k\pmod{p+1}$.
Proof.
The idea behind this proof is similar to that of Lemma 4.12. Recalling that $i_{t-1}=n-k$ and $i_{t-2}\leq n-p-1$ by Lemma 4.11 (c), our immediate goal is to confirm that $i_{t-2}=n-p-1$.
Assume, for the sake of contradiction, that $m_{n-p-1}\notin\sigma$. We claim that $m_{n-p-1}$ cannot be a true gap of $\sigma$. For contradiction, suppose that it is a true gap of $\sigma$. Contradiction is evident if $\sigma$ is absolutely critical because it lacks true gaps. If $\sigma$ is fortunately critical, it is potentially-type-2 and we have $m_{n-p-1}\succ\operatorname{sb}(\sigma)$ by Remark 2.9. Given that $\operatorname{sb}(\sigma)=m_{i_{1}}$ as stated in Lemma 4.10 and the only elements in $\sigma$ dominated by $m_{n-p-1}$ are $m_{n-k}$ and $m_{n}$, we have $\operatorname{sb}(\sigma)=m_{i_{1}}=m_{n-k}$, and therefore $|\sigma|=2$. This yields a contradiction because $\sigma$ cannot have any bridge in this case. Thus, we conclude that $m_{n-p-1}$ cannot be a true gap of $\sigma$. This means adding $m_{n-p-1}$ to $\sigma$ creates a new bridge $m_{b}\in\sigma$ with $m_{n-p-1}\succ m_{b}$. Using Proposition 4.7, there exists a monomial $m_{s}\in\sigma$ such that its distance to $m_{n-p-1}$ along $m_{b}$ is at most $p$. In other words, given that $m_{n-p-1}\succ m_{b}$, we have
  1. (1)
    $s-(n-p-1)\leq p$ when $1\leq n-1-p<b<s\leq n$ or
  2. (2)
    $s+n-(n-p-1)\leq p$ when $1\leq s<n-1-p<b\leq n$.
The latter is untenable since it implies $s<-1$. For the former, $m_{b}=m_{n-k}$ and $m_{s}=m_{n}$, contradicting (1) as the distance between $m_{n-1-p}$ and $m_{n}$ is $p+1$. Hence, $i_{t-2}=n-1-p$. If $i_{t-2}$ equals $i_{1}$, our investigation concludes. Suppose $i_{t-2}\neq i_{1}$.
The next step is to show $i_{t-3}=n-k-(p+1)$. We first claim that $m_{i_{t-2}}$ cannot be a bridge of $\sigma$. If it is a bridge of $\sigma$, then $\operatorname{sb}(\sigma)=m_{i_{1}}$ by Lemma 4.10 because $\sigma$ must be fortunately critical. This means $i_{1}=i_{t-2}$, a contradiction to our assumption. Since $m_{i_{t-2}}$ cannot be a bridge of $\sigma$, we have $i_{t-3}\leq n-k-(p+1)$ by Proposition 4.7. So, it suffices to show $m_{n-k-(p+1)}\in\sigma$.
For the sake of contradiction, suppose $m_{n-k-(p+1)}\notin\sigma$. If $m_{n-k-(p+1)}$ is a true gap of $\sigma$, then $\sigma$ is potentially-type-2 and $m_{n-k-(p+1)}\succ\operatorname{sb}(\sigma)=m_{i_{1}}$ by Remark 2.6 and Lemma 4.10. This means $i_{1}=i_{t-2}$, a contradiction. Since $m_{n-k-(p+1)}$ is not a true gap of $\sigma$, the cell $\sigma\cup m_{n-k-(p+1)}$ has a new bridge $m_{b}\in\sigma$ where $m_{n-k-(p+1)}\succ m_{b}$. Then, by Proposition 4.7, there exists $m_{s}\in\sigma$ such that the distance between $m_{n-k-(p+1)}$ and $m_{s}$ along $m_{b}$ is at most $p$. In other words, given that $m_{n-k-(p+1)}\succ m_{b}$, we have
  1. (1)
    $s-(n-k-(p+1))\leq p$ when $1\leq n-k-(p+1)<b<s\leq n$ or
  2. (2)
    $s+n-(n-k-(p+1))\leq p$ when $1\leq s<n-k-(p+1)<b\leq n$
One can verify that (2) cannot happen. For the first case, the monomial $m_{s}$ is either $m_{n-k}$ or $m_{n}$. For either value of $s$, one can see that the distance is at least $p+1$, a contradiction to (1). Therefore, $i_{t-3}=n-k-(p+1)$. If $i_{t-3}=i_{1}$, the process terminates here. If $i_{t-3}\neq i_{1}$, one can repeat the above arguments for the next steps until reaching $i_{1}=p-k+1$. This results in the congruence $p-k+1\equiv n\text{ or }n-k\pmod{p+1}$, concluding the proof. ∎
Below, we describe all $A_{u}$-critical cells. This proposition serves as the centerpiece of this chapter’s main result.
Proposition 4.14.
Let $n=(p+1)q+r$ where $0\leq r\leq p$. Then, we have:
  1. (1)
    For $r=0$, the only $A_{u}$-critical cells are the cells $\sigma_{i}$, where
    $\sigma_{i}:=\{m_{j}\mid j\equiv i\pmod{p+1}\}$
    for each $1\leq i\leq p$. Moreover, each $\sigma_{i}$ is absolutely critical.
  2. (2)
    For $r\neq 0$, the only $A_{u}$-critical cell is
    $\tau_{r}:=\{m_{j}\mid j\geq r,j\equiv r,2r\pmod{p+1}\}.$
    Moreover, the cell $\tau_{r}$ is absolutely critical if and only if $2r>p$.
Proof.
Let $\sigma=\{m_{i_{1}},\ldots,m_{i_{t}}\}$ be an $A_{u}$-critical cell with $t\geq 2$. To identify all $A_{u}$-critical cells, we first address the immediate case where $n=p+1$. It is evident that $t\geq 3$ is not possible, as this would lead to $\operatorname{sb}(\sigma)=m_{n}$, a contradiction by Lemma 4.11 (a). Consequently, $\sigma$ must be of the form $\sigma_{i}=\{m_{i},m_{n}\}$ for all $1\leq i\leq p$. A straightforward check confirms that every such $\sigma_{i}$ is absolutely critical. Having settled this case, we proceed under the assumption $n>p+1$.
Recall from Lemma 4.12 and Lemma 4.13 that the distance between consecutive elements of $\sigma$ alternates between $k$ and $(p+1)-k$. Specifically, for $1\leq k\leq p$, we have
$\sigma=\{m_{p-k+1},\ldots,m_{n-k-(p+1)},m_{n-(p+1)},m_{n-k},m_{n}\}$
where $p-k+1\equiv n$ or $n-k\pmod{p+1}$. It is important to clarify that the distance between the first and last elements of $\sigma$ is not considered.
Observe that the conditions $p-k+1\equiv n\pmod{p+1}$ and $n-k\equiv 0\pmod{p+1}$ are equivalently expressed as $n\equiv p+1-k\pmod{p+1}$ and $n\equiv 0\pmod{p+1}$, respectively. Consequently, the cell $\sigma$ can be expressed as follows:
$\sigma=\begin{cases}\{m_{(p+1-k)+i(p+1)},m_{(i+1)(p+1)}\leavevmode\nobreak\ :\leavevmode\nobreak\ 0\leq i\leq q-1\}&\text{if }r=0,\\ \{m_{r+i(p+1)},m_{2r+i(p+1)}\leavevmode\nobreak\ :\leavevmode\nobreak\ 0\leq i\leq q-1\}\cup\{m_{n}\}&\text{if }r\neq 0,\end{cases}$
where the former corresponds to $\sigma_{p+1-k}$ for $1\leq k\leq p$ and the latter corresponds to $\tau_{r}$ in the statement of the proposition. It is easily verified that each $\sigma_{p+1-k}$ is absolutely critical, as it lacks both bridges and true gaps for $1\leq k\leq p$. Furthermore, one can verify that $\tau_{r}$ is absolutely critical if and only if $2r>p$. ∎
The conclusions drawn below directly arise from the descriptions of $A_{u}$-critical cells given in Proposition 4.14.
Corollary 4.15.
Let $\sigma\in V(G_{u})$ be a critical cells of ${\mathcal{G}}(I_{p}(C_{n}))$. Then
$|\sigma|=\begin{cases}2q&\text{ if }n\equiv 0\pmod{(p+1)},\\ 2q+1&\text{ otherwise}.\\ \end{cases}$
In particular, all $A_{u}$-critical cells have the same cardinality.
We now present the main theorem of this section, which follows immediately from the preceding corollary, and the description of the differentials of the corresponding generalized Barile-Macchia resolution in Theorem 2.2.
Theorem 4.16.
A generalized Barile-Macchia resolution of $I_{p}(C_{n})$ is minimal.
We conclude this section with a brief discussion. In [9], the first two authors introduced and examined Barile-Macchia resolutions, which are cellular and independent of $𝕜\operatorname{char}(\Bbbk)$. Though effective for many classes of ideals, this method does not always produce a minimal free resolution. Specifically, the edge ideal of a 9-cycle, $I_{2}(C_{9})$, cannot be minimally resolved by a Barile-Macchia resolution, as highlighted in [9]. To the best of our knowledge, no previous construction has been developed to minimally resolve edge ideals of cycles; we introduce the first such construction, ensuring a cellular minimal resolution.
Corollary 4.17.
Edge ideals of cycles have cellular minimal free resolutions.
Echoing the approach of the preceding section, we conclude with an example presenting a CW complex that supports minimal free resolutions of the path ideals of cycles, where the projective dimensions equal $2$. Specifically, these correspond to the ideals $I_{p}(C_{n})$ for which $p+2\leq n\leq 2p+1$.
We consider the minimal generalized Barile-Macchia resolutions of these ideals. One can obtain the following list of all critical cells of $I_{p}(C_{n})$:
,$\emptyset,\{m_{1}\},\dots,\{m_{p+2}\},\{m_{1},m_{2}\},\{m_{2},m_{3}\},\dots,\{m_{n-1},m_{n}\},\{m_{n},m_{1}\},\{m_{r},m_{2r},m_{n}\},$
where $n=(p+1)+r$ for $1\leq r\leq p$. The following $n$-gon $\Delta$ is in correspondence with the critical cell of $I_{p}(C_{n})$:
$\{m_{1},m_{2}\}$$\{m_{2},m_{3}\}$$\{m_{3},m_{4}\}$$\{m_{4},m_{5}\}$$\{m_{1},m_{n}\}$$\{m_{1}\}$$\{m_{2}\}$$\{m_{3}\}$$\{m_{4}\}$$\{m_{5}\}$$\{m_{n}\}$$\{m_{r},m_{2r},m_{n}\}$
Drawing parallels with Example 3.17, and using [6, Lemma 2.2], one can establish that all restricted subcomplexes of $\Delta$ are either empty, a line segment, or the $n$-gon itself, which are all acyclic. Thus, $\Delta$ support the minimal free resolution of $R/I_{p}(C_{n})$.
Acknowledgements. We thank Sara Faridi for helpful discussions about the manuscript. The first author was supported by NSF grants DMS 1801285, 2101671, and 2001368.